OpenStudy (anonymous):

if a^x=b , b^y=c , c^z=a show that xyz=1( abc positive numbers)

OpenStudy (anonymous):

a ^xyz = a so xyz = 1

OpenStudy (anonymous):

x=logb/loga y=logc/logb z=loga/logc

OpenStudy (anonymous):

xyz= logb/loga x logc/logb x loga/logc =1

OpenStudy (anonymous):

how a^xyz=1

OpenStudy (anonymous):

x/(b+c-a) = z/(a+b-c) => (a+b-c)x = (b+c-a)z x/(b+c-a) = y/(c+a-b) => (c+a-b)x = (b+c-a)y So (b+c-a)(z-y) = 2(b-c)x

OpenStudy (anonymous):

a^x=b and b^y = c therefore a^xy = c a^xy = c and c^z = a therefore a^xyz = a now a^1 = a therefore xyz = 1

OpenStudy (anonymous):

thamkz a lot.i got it..

OpenStudy (anonymous):

good

OpenStudy (anonymous):

2^2^x=16^263x solve